5 cards were dealtfrom a deck of 52 cards. Wahts the probability that 4 aces show up?
Answer by stanbon(48568)

5 cards were dealt from a deck of 52 cards. Wahts the probability that 4 aces show up?
--------------
# of ways to pick 4 aces: 4C4 = 1
# of ways to pick the 5th card = 48
----------
# of ways to get a hand of 4 aces and one other card: 1*48 = 48
----------
# of ways to pick 5 cards of any kind 52C5 = 2598960
----------------------------
Prob(5card hand with 4 aces) = 48/2598960 = 0.0000184689...
================
Cheers,
Stan H.